OpenStudy (vera_ewing):

A sample of nickel is heated to 95.0°C and placed in a coffee-cup calorimeter. If the 50.0 g of water in the calorimeter is initially 22.3°C and warms to 25.4°C, what mass of nickel was heated? (The specific heat of nickel is 0.444 J/°C g and water is 4.18 J/°C g). A. 15.4 g B. 57.5 g C. 2.1 g D. 21.0 g

OpenStudy (vera_ewing):

@Somy

OpenStudy (somy):

Q = mc Delta T use m (mass) of water, c (specific heat) of water and temperature change of water Find the Q then use the Q value and the formula again but this time use temperature change of Nickel which is from 95 to 25.4 and use c (specific heat of nickel) so u will have value of Q found earlier u have temp change of nickel and u have specific heat thus find mass using same formula

OpenStudy (vera_ewing):

so what would the answer be?

OpenStudy (somy):

i dont know, calculate it to find out

OpenStudy (vera_ewing):

can you show me how to?

OpenStudy (somy):

Firstly use this formula Q= mc Delta T m= mass of water c= specific heat of water Delta T is temperature change = 25.4°C - 22.3°C = get it and so get the answer for Q

OpenStudy (vera_ewing):

so it's C?

OpenStudy (somy):

did you try calculating and got C?

OpenStudy (vera_ewing):

well i actually got 03.1

OpenStudy (vera_ewing):

which is none of the answer choices lol

OpenStudy (somy):

you haven't finished tho

OpenStudy (somy):

+ how can u exactly get that can u show me what u did?

OpenStudy (vera_ewing):

i just did 25.4-23.3

OpenStudy (somy):

lol thats just temperature = Delta T you havent used the formula

OpenStudy (vera_ewing):

oh wait so then what do i do?

OpenStudy (somy):

lel try reading what i said earlier :3

OpenStudy (somy):

m= mass of water 50g c= specific heat of water = 4.18 J/°C g Delta T is temperature change = 25.4°C - 22.3°C = 3.1°C so use all the values and put them into the formula Q= mc Delta T

OpenStudy (vera_ewing):

hmm ok well i'm not really sure how to do it. can you just do it and explain the steps to me?

OpenStudy (somy):

and you will get the value for Q and we are not done there yet

OpenStudy (somy):

\[\sf Q = m \times c \times \Delta T\] \(\sf \Delta T\), that triangle is sign of Delta and means 'change', in other words change in temperature

OpenStudy (somy):

i gave u all the values above so put them into the formula its just multiplication, im sure you can do it :3

OpenStudy (somy):

write the value you get then we'll move on to second step